### v1

# SemNav: A Semantic Segmentation-Driven Approach to Visual Semantic Navigation

Abstract
Visual Semantic Navigation (VSN) is a fundamental problem in robotics, where an agent must navigate toward a target object in an unknown environment, mainly using visual information. Most state-of-the-art VSN models are trained in simulation environments, where rendered scenes of the real world are used, at best. These approaches typically rely on raw RGB data from the virtual scenes, which limits their ability to generalize to real-world environments due to domain adaptation issues. To tackle this problem, in this work, we propose SemNav, a novel approach that leverages semantic segmentation as the main visual input representation of the environment to enhance the agent’s perception and decision-making capabilities. By explicitly incorporating high-level semantic information, our model learns robust navigation policies that improve generalization across unseen environments, both in simulated and real world settings. We also introduce a newly curated dataset, i.e. the SemNav dataset, designed for training semantic segmentation-aware navigation models like SemNav. Our approach is evaluated extensively in both simulated environments and with real-world robotic platforms. Experimental results demonstrate that SemNav outperforms existing state-of-the-art VSN models, achieving higher success rates in the Habitat 2.0 simulation environment, using the HM3D dataset. Furthermore, our real-world experiments highlight the effectiveness of semantic segmentation in mitigating the sim-to-real gap, making our model a promising solution for practical VSN-based robotic applications. We release SemNav dataset, code and trained models at: https://github.com/gramuah/semnav
I Introduction
Autonomous navigation remains a fundamental challenge in robotics, particularly in unstructured and dynamic environments. Traditional approaches, such as the ones using Simultaneous Localization and Mapping (SLAM), for example, focus on a geometric reconstruction of the environment in conjunction with path planning and obstacle avoidance techniques (e.g. [1, 2, 3]). However, these classical models typically struggle when they need to generalize across environments and domains [4].
Visual Semantic Navigation (VSN) has emerged as an alternative paradigm that leverages advances in machine learning and the availability of large-scale data. VSN models are essentially learning-based navigation approaches that apply end-to-end methods to directly map sensory visual inputs to actions that control the robot (e.g. [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]). VSN approaches, instead of planning a geometric route over a map, learn navigation policies that enable robots to explore unknown environments while following semantically meaningful instructions, such as “Move towards a chair” or “Find a bedroom”. This particular problem is known as the Object Navigation (ObjectNav) problem [16], where an agent must navigate, in an unknown scenario, so without any map, toward an instance of a given object category using mainly visual information. All these VSN ObjectNav systems are trained and evaluated in complex simulation platforms, such as Habitat [17, 16], AI2-THOR [18] or ProcTHOR [19], where the agents learn to navigate in interactive 3D environments and complex physics-enabled scenarios. AI2-THOR and ProcTHOR enable interaction with realistic virtual 3D environments generated through computer graphics. Habitat, on the other hand, allows agents to interact with 3D models corresponding to real-world 3D scans of scenes and houses, such as those included in the HM3D datasets [20, 21].
Figure 1: Traditional VSN models are trained in simulation environments where they primarily rely on RGB images. However, due to the significant domain gap between the real world and simulated images, VSN systems often fail to navigate when deployed on real robots. To address this issue, we propose the SemNav paradigm, in which VSN systems use semantic segmentation of the environment as their primary input sensor.
Despite recent progress, state-of-the-art ObjectNav models often struggle with real-world generalization due to the significant domain gap between simulation-trained policies and real-world execution [22, 23, 24]. In other words, no matter how realistic the simulation environments used during learning are, the real world presents other challenges. To begin with, the images provided by the simulation platforms are either not entirely realistic or exhibit artifacts typical of 3D scanning processes of real-world scenes. The egocentric images that a robotic platform will acquire in the real world will present higher levels of blurring due to the platform’s movement, more diverse lighting situations, and will also include changes that naturally occur in dynamic environments. This results in navigation strategies that fail to generalize beyond the training domain. To address this challenge, we propose SemNav (see figure 1), a novel ObjectNav approach that uses semantic segmentation as the main visual input. Our goal in basing the visual perception of the VSN system on semantic segmentation is twofold. First, we argue that a representation based on semantic segmentation naturally mitigates the domain adaptation problem between the real world and simulated environments. The differences between the semantic segmentations of a real-world scene and those of a rendered scene in a simulation environment such as Habitat [17] are smaller than those found in RGB images from both domains. Second, by incorporating semantic priors into the visual representation of the environment, our model learns robust navigation policies that better adapt to unseen environments, hence improving the success of the navigation task. Our approach is inspired by the intuition that human navigation heavily relies on semantic cues—understanding that a “kitchen” typically contains a “sink” or a “refrigerator” enables more efficient exploration.
To validate our approach, we conduct extensive experiments in both simulated and real-world environments. Our findings indicate that semantic segmentation significantly improves navigation performance, particularly in bridging the sim-to-real gap. Unlike conventional methods that degrade in real-world deployments, SemNav exhibits higher robustness by leveraging structured semantic information. Overall, our main contributions in this work are as follows:
  • We release the new SemNav dataset designed for training semantic segmentation-aware navigation models, enabling further research in this domain.
  • We present SemNav, a novel ObjectNav model designed to employ the semantic segmentation of the robot’s egocentric view for learning the mapping between visual observations and navigation actions in unexplored environments. SemNav is versatile, demonstrating robust and efficient navigation and exploration behaviors in both simulated environments and real-world scenarios.
  • In our experimental evaluation, the SemNav approach demonstrates superior performance compared to state-of-the-art methods, both on the Habitat 2.0 simulation platform [16] and in real-world testing.
II Related work
Robotic autonomous navigation has long been a key focus in research, with many methods developed to tackle the challenge of maneuvering robots through complex and dynamic environments. Classical navigation approaches are mainly based on Simultaneous Localization and Mapping (SLAM) [1, 2, 3, 25, 26, 27]. SLAM-based solutions primarily tackle the navigation problem by focusing on obstacle avoidance, map creation and path-planning algorithms. However, these models face challenges in generalizing to different environments [4].
To address these limitations and capitalize on recent breakthroughs in machine learning and the availability of large-scale datasets, research has increasingly shifted toward learning-based approaches for robotic navigation. Within this category, Visual Semantic Navigation (VSN) models stand out. Unlike traditional methods that depend on geometric path planning over pre-constructed maps, VSN models adopt a fundamentally different strategy by learning navigation policies that enable robots to navigate unfamiliar environments based on semantically meaningful instructions, such as “Go to a chair” (ObjectNav problem) or “Find this image” (ImageNav task). For a comprehensive overview of recent progress and challenges in VSN, interested readers may refer to [16]. Fundamentally, VSN models learn to make navigation decisions based mainly on visual information (RGB and/or depth data) [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Two main VSN training approaches include both Imitation Learning (IL) and Reinforcement Learning (RL). IL-based methods learn navigation policies from previously annotated demonstrations [5, 6]. RL-based methods can be split into two categories. The first category consists of models that follow an end-to-end RL approach [28, 9, 10, 11] via interaction with the environment. Several training strategies have been proposed for these systems, including the use of auxiliary tasks [12], object-relationship graphs [13], and the fusion of visual and auditory information [29, 30]. The second category includes works that employ modular learning approaches [8, 22, 31, 23], where the navigation process is divided into independent modules addressing different aspects of navigation.
In recent years, advances in Large Language Models (LLMs) have enabled their application to the visual semantic navigation problem, giving rise to Vision-Language Navigation (VLN) models [32, 33, 34]. In these models, an embodied agent executes natural language instructions within real 3D environments to navigate. Finally, we have the recent diffusion-based navigation approach where the latest advancements in generative modeling, particularly in diffusion models [35, 36] have influenced robot navigation [37, 38].
As previously argued, despite significant efforts in developing VSN models, a major limitation persists: most models are trained and evaluated predominantly in simulation environments. In fact, only a few studies have attempted to evaluate VSN solutions in real-world settings [22, 23, 24]. The results reported in these studies reveal a significant domain gap problem: while VSN models achieve high success rates in simulation, their performance deteriorates drastically when deployed in real-world environments. In this work, we propose a novel VSN model, named SemNav, designed to learn navigation policies for the ObjectNav problem using a semantic segmentation of the robot’s egocentric view. The semantic segmentation output is fed into a CNN-based visual encoder, which extracts navigational cues from the segmentation information to enhance navigation in unseen environments. To the best of our knowledge, no previous VSN work has implemented this idea as we have. Some studies leverage semantic segmentation to define navigable regions and plan routes [39, 40], while others combine this information with object detection systems [41] to enable target-driven visual navigation. Our initial hypothesis, confirmed by experimental results, is that semantic segmentation serves as a valuable source of information that allows: (1) the development of more efficient VSN models and (2) the natural mitigation of domain adaptation issues.
III SemNav Approach
III-A SemNav dataset
The first essential requirement for training a VSN model that utilizes semantic segmentation as its visual input is the availability of a dataset containing this type of information. In this work, we have built upon the HM3D dataset [5] to construct the SemNav dataset. Specifically, in this section, we detail the process through which we enriched HM3D to enable the generation of a semantic segmentation sensor compatible with the Habitat simulator [17]. This semantic sensor allows any agent interacting with Habitat to access an egocentric view that corresponds to the semantic segmentation of the scene currently perceived by the robot.
The HM3D dataset provides a collection of 3D real-world spaces, densely annotated with semantic information, as illustrated in Figure 2. This dataset comprises over 140,000 object instance annotations distributed across 216 3D environments and approximately 3,100 rooms. The semantic data is embedded in the form of texture images mapped onto the original 3D geometry of the HM3D scenes. This information is packed into binary glTF format files, one per scene. Each object instance is assigned a unique color identifier, which maps to a corresponding textual label specifying the object’s category, stored in a separate annotation file.
The separation of object texture information and semantic labels into different files leads to a limitation: the built-in sensor in Habitat for the HM3D dataset does not deliver a true semantic segmentation of the scene. For instance, chairs, as illustrated in Figure 2, are not consistently assigned the same identifier or label across different scenes.
Our model, SemNav, relies on semantic segmentation to navigate effectively. To meet this requirement, we have developed an enhanced semantic segmentation sensor. It builds upon the original semantic information available in the HM3D dataset. Unlike the default information, our sensor ensures that object categories are consistently assigned unique and uniform labels across all scenes. Technically, the creation of these sensors has been automated by designing a mapping process that allows translating the information associated with textures into semantic segmentation labels that are now expanded both intra- and inter-scenes. This improvement guarantees that different instances of the same object category receive identical labels within a single scene and across multiple scenes, as illustrated in Figure 2.
To provide flexibility and address varying levels of semantic granularity, we have created two semantic segmentation sensors: one with 1,630 distinct object categories, referred to as SemNav 1630, and another with a more compact set of 40 categories, called SemNav 40. The SemNav 1630 sensor was generated by leveraging the fine-grained object annotations available in HM3D [5], where each manually labeled object instance was assigned to a unique category. This allowed identically annotated objects to be grouped under the same category. However, upon conducting a detailed analysis, we identified certain limitations in these annotations, including noise from overly specific object labels and occasional misclassifications. To address these challenges and improve the robustness of the semantic segmentation, we created a second semantic segmentation sensor, SemNav 40. In this version, each of the 1,630 original annotations was manually mapped to one of the 40 broader categories defined by the NYUv2 dataset [42], a widely recognized benchmark for indoor semantic segmentation. This mapping reduced the initial set of fine-grained categories to a more semantically coherent set of 40, thereby simplifying the navigation process and minimizing the impact of noisy annotations. Both SemNav 1630 and SemNav 40 were integrated into the Habitat simulator, allowing any agent to query the simulator for true semantic segmentation corresponding to its egocentric view in the 3D environment being rendered.
From this point onward, we will refer to each sensor as if it were a dataset in its own right. Accordingly, we have released two datasets, SemNav 1630 and SemNav 40, which are fully integrated into the Habitat platform. This integration enables straightforward training of any model requiring semantic segmentation input, streamlining the development of advanced navigation models and reducing implementation overhead. Both datasets are accessible at https://github.com/gramuah/semnav.
Figure 2: Comparison between the HM3D dataset and the SemNav 1630 and SemNav 40 datasets. In HM3D, objects like chairs are inconsistently labeled with different colors across and within scenes. In contrast, SemNav 1630 assigns a uniform color to objects of the same category while reflecting finer distinctions, such as differentiating kitchen and dining tables. SemNav 40 further simplifies the labeling by merging less critical categories, grouping both types of tables under a shared label to streamline navigation tasks. Best viewed in colour.
III-B SemNav model
In this work, we also introduce the SemNav model, designed to address the VSN problem known as ObjectNav [16]. The objective of an ObjectNav navigation episode is to enable an agent to navigate in a scene $S_{i}$ from a set of available scenes $\mathcal{S}=\{S_{1},\dots,S_{n}\}$, towards an object of a specific category $c_{i}$ belonging to the category set $\mathcal{C}=\{c_{1},\dots,c_{m}\}$, starting from an initial position $p_{0}$ in the navigation environment.
For our SemNav model, we define the navigation task as follows. Given a target object class $c_{i}$, e.g., chair, our goal is to navigate to an instance of this category using only the semantic segmentation of the environment. Figure 3 shows the architecture of the proposed SemNav model. We define a VSN model in which, at each position $p_{i}$, the agent has access to an environmental observation $o_{p_{i}}$, represented as the tuple:
$o_{p_{i}}=\{o^{\text{semantic}}_{i},o^{\text{compass}}_{i},o^{\text{gps}}_{i}\}\;.$
As shown in Figure 3, the SemNav model receives as inputs: the semantic segmentation of the environment ($o^{\text{semantic}}_{i}$); the relative orientation with respect to the initial viewpoint ($o^{\text{compass}}_{i}=\Delta\alpha_{i}$); and the relative displacement from the starting position ($o^{\text{gps}}_{i}=(\Delta x_{i},\Delta y_{i},\Delta z_{i})$).
Figure 3: Proposed architecture for the SemNav model.
[…]
For the SemNav model, we adopt an IL strategy based on human demonstrations to learn the parameters in $\theta$. Specifically, in our experiments, we leverage a dataset of 77k human-driven navigation trajectories from HM3D [5], which are also applicable to the novel SemNav dataset. A trajectory $\tau$ of length $T$, representing a human demonstration, is defined as a sequence of observation-action tuples:$\tau=(o_{p_{0}},a_{p_{0}},o_{p_{1}},a_{p_{1}},\dots,o_{p_{T}},a_{p_{T}})$. The set of trajectories used for training is denoted as $\mathcal{T}=\big{\{}\tau_{i}\big{\}}_{i=1}^{N}$. Following an IL approach, we optimize the parameters $\theta^{*}$ of our navigation policy by solving the following optimization problem
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\sum_{i=1}^{N}\sum_{(o_{p_{t}},a_{p_{t}})\in\tau_{i}}^{T}-\log(\pi^{\textsc{SemNav}}_{\theta}(a_{p_{t}}\mid o_{p_{t}}))\;,$ (1)
where the navigation policy is trained by minimizing the discrepancy between the probability distribution for the actions assigned by the model to each observation $o_{p_{t}}$, and the actions of the human experts.
The architecture depicted in Figure 3 comprises several key components. To process the semantic segmentation input, we first transform it into an RGB format. Each pixel is assigned a semantic segmentation label, which is then mapped to a specific color in the RGB space using the semantic segmentation sensor integrated into the SemNav dataset. The encoded input is subsequently processed by a ResNet50 architecture [43], which extracts high-level visual features for downstream tasks. In our experiments, we evaluate two types of initialization for the ResNet50: (1) a random initialization and (2) an initialization using DINO [44]. The input corresponding to the target object category is converted into a one-hot vector and processed through an embedding layer. GPS and compass observations are directly passed through a fully connected layer. All these processed inputs are concatenated and subsequently fed into a Gated Recurrent Unit (GRU) network [45]. This recurrent network is responsible for predicting the discrete action distribution that forms the output of our model.
IV Experiments
IV-A Experimental settings and evaluation metrics
Our main goal in this work is to develop VSN solutions capable of navigating efficiently both in the virtual environments where they are trained and in the real world. For this purpose, we have designed two types of experiments.
Experiments in the simulation environment
For this experimental evaluation, we addressed the ObjectNav task detailed in the Habitat 2023 challenge [16]. We used the official HM3D [21] training and validation sets. For our SemNav model we employed the two semantic segmentation sensors detailed in Section III-A. For evaluation metrics, we followed the standard proposals for the ObjectNav problem [16]. Specifically, these metrics include: 1) the Success Rate (SR), defined as the percentage of navigation episodes successfully completed by the model; 2) the Shortest Path Length (SPL), which compares the distance traveled by the agent during the episode with the shortest possible path length to the target object; 3) the average number of collisions the agents experience with the environment (C); and 4) the average distance to the goal (DTG), which is the mean distance of the agent from the nearest target object at the end of a navigation episode.
Experiments in the real world
For real-world experiments, the evaluations were conducted in a domestic environment, whose floor plan is shown in Figure 4. We employed a TurtleBot 2 robotic platform [46] specifically adapted for the ObjectNav problem. This adaptation replicates the agent’s characteristics during training. Among the modifications, we added a mast to the TurtleBot 2, raising the camera to 1.25 m to match the simulation setup. The camera used is an Orbbec Astra with depth perception. Since our SemNav model requires a semantic segmentation sensor, for real-world testing, we integrated the ESANet model [47], pre-trained on the NYUv2 dataset [42]. This model takes the robot’s egocentric vision as input in RGB+Depth format and generates a semantic segmentation image, which is then fed into our SemNav model.
The embedding of our SemNav model and the state-of-the-art PirlNav [48] in the robotic platform was done using the ROS4VSN library [24], which allows for the integration of VSN solutions into real robots using ROS. We conducted navigation experiments toward five specific object classes: chair, bed, toilet, television monitor, and sofa. For each episode, an object category was selected as the navigation target, and the navigation started from a predefined location and orientation. These starting points were chosen to ensure that: 1) there was no direct visibility of the target object; and 2) there was a minimum distance of 10 meters to the object. The evaluation metric used was the SR, where a navigation episode is considered successful if the robot infers the stop action while being within one meter of the target object. A failure is recorded if the robot collides and cannot proceed or if it requires more than 225 actions to complete the task.
Calculating the SPL metric in the real world is challenging, so we introduce a new metric, Success Divided by Step (SDS), as an alternative for evaluating navigation efficiency in real environments. SDS is computed as $SDS=\frac{N_{TS}}{N_{AS}}$, where $N_{TS}$ is the number of successful trajectories, and $N_{AS}$ is the total actions taken during those successful trajectories to reach the target object category.
Figure 4: Top-down view of the house where ObjectNav experiments were conducted for five object categories.
We release all code for training and evaluating our models, both in simulated environments and in the real world, in the following anonymous repository.
IV-B Results in simulation
IV-B1 Ablation study
In this analysis, we aim to present the impact of the following aspects on the performance of our SemNav model: 1) We test a SemNav configuration that uses only semantic segmentation as visual input—SemNav-Only Semantic Segmentation (SemNav-OS); 2) We add a standard RGB visual input to the SemNav model, which passes through its corresponding visual encoder (a ResNet 50) to be concatenated with the rest of input features—SemNav-RGB + Semantic Segmentation (SemNav-RGBS); 3) Each of these configurations is tested with semantic segmentation using 1630 and 40 categories, as provided in the SemNav dataset; and 4) For each test, we train the model either from randomly initialized weights or by pre-initializing them using DINO [44], since this initialization strategy has demonstrated a significant impact on previous VSN models, e.g., PirlNav [48]. In total, we propose an evaluation of up to eight different configurations for our SemNav model.
Categories dataset Model SR (↑) SPL (↑) C (↓) DTG (↓)
SemNav 1630 OS 65.1 0.29 57.11 2.99
OS+DINO 68.4 0.30 50.39 2.75
RGBS 69.2 0.31 50.59 2.73
RGBS+DINO 70 0.32 39.17 2.69
SemNav 40 OS 72.9 0.34 43.25 2.55
OS+DINO 73 0.34 48.49 2.61
RGBS 74.3 0.35 40.70 2.50
RGBS+DINO 76.2 0.36 38.50 2.36
TABLE I: Evaluation of the performance of different configurations of our SemNav model. We report the metrics: Succes Rate (SR), Shortest Path Length (SPL), average number of collisions (C), and average distance to the goal (DTG).
Table I presents the performance of all these configurations on the SemNav-dataset validation set. In light of the results obtained in this study, our model SemNav +RGBS has provided the best performance. It is interesting to highlight that all models trained with 40-category semantic segmentations have outperformed those trained with 1630-category semantic segmentations. We believe this is primarily due to the annotation issues present in the 1630-category set, as detailed in Section III-A, and the fact that the semantic information necessary for successful navigation is already included in the 40 categories, making the 1630-category segmentation unnecessarily complex. Lastly, our SemNav model does not appear to be significantly affected by initialization with DINO [44], unlike other state-of-the-art models. In our case, DINO initialization slightly improves the results.
This ablation study also demonstrates that semantic segmentation is the key factor impacting model performance. Comparing any of our models with and without RGB input—i.e., OS vs. RGBS—further confirms this observation. For instance, in terms of SR, for the SemNav 40 dataset, the SemNav model improves from 72.9 to 74.3 when using the OS and RGBS versions, respectively. We conclude that these results validate the suitability of the SemNav approach for the VSN problem. Adding RGB information to the model consistently enhances its performance, and for real-world navigation, this would be the optimal architecture to embed in a robotic platform, as we will demonstrate in Section IV-C.
IV-B2 Comparison with the state of the art
We compare here the performance of our SemNav models with state-of-the-art approaches for the ObjectNav problem. We have used the validation set of HM3D [21] dataset. All compared methods use the same experimental evaluation protocol for ObjectNav [16]. The only difference in terms of training data is that our SemNav models use the semantic segmentation information contained in the SemNav dataset.
Model SR (↑) SPL (↑)
DD-PPO [49] 27.9 14.2
OVRL-v2 [50] 64.7 28.1
RRR [51] 30.0 14.0
Frontier based exploration [52, 22] 26.0 15.2
Habitat-Web [5] 57.6 23.8
PirlNav (only IL) [48] (our impl.) 60.9 0.26
PirlNav (only IL) [48] 64.1 0.27
[…]
XGX [23] 72.9 0.36
SemNav-OS 72.9 0.34
SemNav-OS+DINO 73 0.34
SemNav-RGBS 74.3 0.35
SemNav-RGBS+DINO 76.2 0.36
SemNav-RGBS+DINO $\rightarrow$ RL 77.75 0.40
TABLE II: Comparison of SemNav approaches with the state-of-the-art models in the ObjectNav task of the HM3D dataset (validation set). We report the performance using the metrics: Succes Rate (SR) and Shortest Path Length (SPL).
Table II includes, to the best of our knowledge, a comparison with the models that define the state of the art in the ObjectNav problem. Based on the results, we would like to highlight the following conclusions. First, it is noteworthy that our SemNav-OS model, the simplest one as it only uses semantic segmentation information from the environment, outperforms all state-of-the-art models except XGX [23]. We believe this underscores the importance of semantic segmentation information for navigation tasks. Second, when incorporating RGB information as a sensor, our SemNav-RGBS model surpasses all other approaches to the problem, setting a new state of the art on the validation set of the HM3D dataset. Third, if we apply finetuning to our best model using RL (SemNav-RGBS+DINO $\rightarrow$ RL), following the strategy in [48], we achieve an additional improvement, bringing the SR to a value above 77%. And fourth, in the qualitative results, we observed an interesting exploratory capability in the SemNav models. Egocentric videos recorded during evaluations in the simulation environment show how SemNav models tend to explore the environment, probing the different rooms they encounter. If a room is not useful for reaching the target object category, they proceed to leave it and continue exploring. The videos show that SemNav models can traverse hallways, inspect rooms, and move between different floors of a building. Moreover, they do not exhibit erratic movements during navigation, reflecting efficient and consistent decision-making. The promising results obtained in the evaluation metrics support these qualitative observations. It is worth mentioning that these behaviors have also been observed in experiments conducted in the real world (see Section IV-C).
Figure 5 illustrates some of these qualitative navigation results. For example, Figure 5 (a) shows how the agent begins its search in a living room, where it does not find the desired object—in this case, a television screen. It then explores all the rooms on the current floor. After failing to locate the object, the agent moves to the upper floor, where it finally identifies the requested object. Figure 5 (b) depicts another interesting navigation episode. Initially, the agent is in a large living room where it cannot detect the requested object category. In the following images, the agent moves through the environment, passing through two bedrooms and viewing the living room from a different perspective. Finally, the agent detects a television and executes a stop action next to it.
Figure 5: Qualitative results in simulated environments. From top to bottom, these figures show the trajectory followed by the agent in the environment. On the left, the model’s input is illustrated as an RGB image; in the center, the semantic segmentation input perceived by the SemNav model; and on the right, the map reflecting the agent’s path.
IV-C Results in the real world
[…]
SR / Actions SDS
Chair Bed Toilet Sofa TV Monitor Average
PirlNav [48] 0% / 66 0% / 81 0% / 82 0% / 225 0% / 74 0% / 105.6 0
SemNav-OS 100% / 36 0% / 164 0% / 55 100% / 35 100% / 86 60% / 75.2 0.019
SemNav-RGBS 100% / 40 0% / 71 0% / 140 100% / 49 100% / 94 60% / 78.8 0.016
[…]
TABLE III: SR and number of actions per object category, and SDS, reported in real-world experiments.
Figure 6: Comparison of the SR reported in the real world and the simulation environment for the different VSN architectures used.
None of the models were able to reach the bed object, as it was not present in the navigable scene. Interestingly, our SemNav models attempted to climb the stairs when tasked with finding the bed. This behavior arises because, in the simulated environment, the agents can navigate stairs. Moreover, beds are typically located on upper floors, which highlights the type of behavior encoded in the learned navigation policy. Unfortunately, in the real world, our robot lacks this capability. Figure 7 presents some qualitative results.
Figure 7: Qualitative results of the robot successfully navigating in the real world toward a sofa, a television, and a chair.
We conclude that although our SemNav models did not always successfully reach the target category, their navigation exhibited a clear intention to move toward it. The robotic platform’s inability to traverse stairs penalized our models. Finally, in real-world experiments, the SemNav models demonstrated superior performance compared to the other state-of-the-art models. This suggests that semantic segmentation helps mitigate the domain gap between real and simulated environments, enabling more accurate navigation in real-world scenarios too.
V Conclusion
In this paper, we introduced SemNav, a novel VSN model that integrates semantic segmentation as the main visual input to improve navigation efficiency and generalization in unknown environments. Unlike conventional VSN methods that rely on raw RGB images and struggle with sim-to-real transfer, SemNav benefits from the structured semantic segmentations representations, enabling more robust decision-making in both simulated and real-world environments. To support our model, we have released the SemNav dataset, designed for training semantic segmentation-aware VSN models, enabling further research in this direction.
Our extensive evaluation shows that SemNav outperforms state-of-the-art ObjectNav models in both the Habitat 2.0 simulator and real-world tests. By leveraging semantic segmentation priors, our model achieves higher success rates and shorter path lengths, even in unseen environments, and exhibits enhanced adaptability in real-world scenarios. The integration of semantic segmentation also reduces the domain gap, a persistent challenge in VSN research.
Despite SemNav ’s strong performance in simulation and real-world scenarios, its reliance on semantic segmentation data may limit its applicability in environments where pre-trained models are ineffective, such as forests or underwater settings, due to dataset scarcity and labeling challenges. Expanding its applicability to other domains may require labor-intensive annotation efforts, as getting additional semantic segmentation datasets remains a significant challenge, although weakly-supervised strategies could be explored.